OpenStudy (anonymous):

find the area: y=cos(3x)cosx given points ((pi/6),0))

OpenStudy (anonymous):

\[\int\limits_{0}^{pi/6}\cos(3x)\cos(x)\]

OpenStudy (anonymous):

\[1/2\int\limits_{0}^{pi/6}\cos(2x)+\cos(4x)\]

OpenStudy (anonymous):

\[1/2[(\sin(2x)/2) + (\sin(4x)/4)\]